• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Makis Petratos

    Log entry time 00:00:56 on December16,2007

    Entry number 221274

    keyword=Shift Summary

    16:05 Set Left Collimator OPEN (0.8708)

    16:10 Started setting Right HRS at 0.74984 GeV (+2%)

    16:20 Start Run 4043 Left HRS Coll. OPEN 7-foil-opt-tgt
    16:25 End Run 4043

    Alex started doing in parallel NAI HV scans (see separate note)

    16:33 Start Run 21963 Right HRS SIEVE 1.5M tgt:opt
    16:55 End Run 21963

    16:57 Start Run 21964 Right HRS SIEVE 0.5M tgt:opt
    17:05 End Run 21964

    17:10 Set Right HRS Coll OPEN

    17:13 Start Run 21965 Right HRS Coll OPEN 1,4M tgt:opt
    17:18 End Run 21965

    Started setting Left HRS at 0.74984 GeV

    17:22 Start Run 4050 Left HRS Coll OPEN tgt:opt 1.7M
    17:28 End Run 4050

    17:32 Start Run 21968 Right HRS Coll OPEN tgt:opt 1.5M
    17:38

    17:45 Set Left HRS Collimator SIEVE

    17:43 Start Run 4051 Left HRS SIEVE tgt;opt 1.5M
    18:05 End Run 4051

    18:10 Set Left HRS Coll OPEN

    18:10 Started Setting Right HRS to 0% (E=.73514 GeV)

    18:35 RF problem, down for 15 mins?

    18:45 Beam is back

    18:51 Start Run 21967 Right Coll OPEN tgt:opt 1.M
    18:55 End Run 21967

    19:00 Set Right HRS Collimator SIEVE

    19:01 Setting Left HRS to 0% - Set Left Collimator SIEVE

    19:03 Start Run 21968 Right HRS Coll SIEVE tgt:opt

    19:09 Start Run 4057 Left HRS Coll SIEVE tgt:opt

    19:20 End Run 21968

    19:26 Right HRS Collomator OPEN

    19:42 Left HRS Collimator OPEN

    19:44 Start Run 4058 Left HRS Coll OPEN tgt:opt
    19:51 End Run 4058

    _________________________________________
    19:55 ==== Moving to Dummy-10cm tgt ====
    -----------------------------------------

    ----- Runs 4059 & 4060 (Left HRS) --> JUNK

    20:12 Start Run 4061 (Left HRS) Dummy-10cm
    20:13 Start Run 21969 (Right HRS) Dummy-10cm

    20:22 End for Runs 4061 (Left) & 21969 (Right)

    _________________________________________
    20:23 ==== Moving to Dummy-15cm tgt ====
    -----------------------------------------

    20:28 Start Run 4062 (Left HRS) Dummy-15cm
    20:28 Start Run 21970 (Right HRS) Dummy-15cm

    20:37 End for Run 21970 (Right)
    20:41 End for Run 4062 (Left)

    ________________________________
    20:42 ==== Moving to C-tgt ====
    --------------------------------

    20:54 Start Run 4063 (Left) C-tgt
    20:55 Start Run 21971 (Right) C-tgt

    21:01 End of Runs 4063 (Left) & 21971 (Right)

    ________________________________
    21:02 ==== Moving to Fe-tgt ====
    --------------------------------

    21:07 Start Run 4064 (Left) Fe-tgt (@ 50uA)
    21:08 Start Run 21972 (Right) Fe-tgt (@ 50uA)

    21:14 End of Runs 4064 & 21972

    21:17 Start Run 4065 (Left) Fe-tgt (@ 10uA)
    21:18 Start Run 21973 (Right) Fe-tgt (@ 10uA)

    21:27 End of Runs 4065 & 21973

    ________________________________
    21:30 ==== Moving to Pb-tgt ====
    --------------------------------

    21:41 Start Run 4066 (Left) Pb-tgt (@ 5 uA)
    21:42 Start Run 21974 (Right) Pb-tgt (@ 5 uA)


    21:51 End of Runs 4066 & 21974

    ________________________________
    21:53 ==== Moving to He-tgt ====
    --------------------------------

    22:07 Start Run 4067 (Left) He-tgt
    22:08 Start Run 21975 (Right) He-tgt

    22:14 End of Runs 4067 & 21975


    _________________________________________________________________
    22:15 ==== Moving to Optics target for HV scan for NaI det.
    -----------------------------------------------------------------
    22:16 ==== Changing the Right HRS momentum to 720.433 MeV/c (-2%)
    -----------------------------------------------------------------

    22:27 Alex started HV scans for NaI

    23:10 Start Run 21976 (Right) optics-tgt & open Collim. @ -2%
    23:17 End of Run 21976

    23:24 Sieve slit IN at Right HRS
    23:25 HV NaI scan completed with Left/HRS --> Setting L/HRS to -2%
    23:25 Sieve slit IN at Left HRS

    23:33 Start Run 4074 (Left) at -2% with sieve slit IN & optics-tgt
    23:35 Start Run 21977 (Right) at -2% with sieve slit IN & optics-tgt

    00:00 Runs 4074 & 21977 Ended